OpenStudy (anonymous):

What is the power of a crane that lifts 500kg 3 metres in 2 seconds?

OpenStudy (anonymous):

power of crane=workdone by crane/required time =change in potential energy of the block/required time =mgh-0/2 =500*9.8*3/2=14700/2=7350watt=9.8525 Hp